# PhysGaia: A Physics-Aware Dataset of Multi-Body Interactions for Dynamic Novel View Synthesis

### v1

@@ 3.1 Dataset Properties
In contrast to real-world video datasets, which are easier to capture but lack access to underlying physical properties, our simulated dataset provides complete ground-truth physical information. This includes 3D particle trajectories and accurate physics parameters such as viscosity, Young’s modulus, Poisson’s ratio, and temperature for gas-based scenes [^3] Temperature is a crucial component for buoyancy calculation in smoke simulations.. This comprehensive provision enables precise evaluation of physical reasoning in dynamic scenes and directly supports the future research directions discussed in Section 3.2.
@@
Our dataset uniquely supports both multiview and monocular DyNVS task. Unlike most multiview datasets [2, 50, 51, 52] with videos captured from fixed camera positions like CCTV, PhysGaia provides several moving monocular video sequences with independent trajectories as training data. This allows for diverse training configurations; using full sequences corresponds to a multiview DyNVS task, while using a single sequence corresponds to a monocular DyNVS task. The monocular setup matches datasets like DyCheck [13], providing realistic handheld camera inputs without synthetic teleportation. For evaluation, we employ two static cameras with a large baseline.
@@
Figure 4: Examples of diverse modalities that users can generate from the provided simulation node graphs.
We provide the complete simulation node graphs and their exact parameter settings used to construct our dataset. These graphs encompass all relevant components, including physics solvers, source geometries, camera positions, lights, materials, and texture controls. By modifying these nodes, users can easily generate customized scenes or additional modalities such as depth maps, surface normals, and re-lighted images tailored to their specific downstream tasks, as shown in Figure 4.
@@
Our PhysGaia benchmark is designed for research-friendly and easy access, providing integration pipelines that enable the use of state-of-the-art DyNVS models [15, 36, 37, 38] with our data, as detailed on our GitHub repository [^4] https://github.com/mjmjeong/PhysGaia. Furthermore, we include COLMAP-reconstructed point clouds for each scene. These provisions aim to foster the adoption of our dataset by researchers working for state-of-the-art DyNVS models.
@@ 3.2 Potential Research
We highlight the potential impact of our PhysGaia dataset by outlining several promising future research directions it uniquely enables.
Advancing physical reasoning in dynamic scenes
@@
Since our PhysGaia dataset provides ground-truth physics information, it facilitates precise evaluation of physical reasoning in dynamic scenes. For instance, ground-truth physics parameters like viscosity can be used to evaluate inverse physics estimation methods, where differentiable simulators are employed to optimize these parameters. Furthermore, unlike existing DyNVS research that primarily focuses on photorealism and relies solely on ground-truth RGB images, our dataset offers ground-truth 3D trajectories. This enables evaluating the actual motion of individual Gaussian primitives in recent GS-based methods, moving beyond just visual fidelity. We believe this unique feature establishes our dataset as a valuable benchmark for developing and evaluating physics-aware DyNVS models./asset
Exploring multi-body interactions
@@
While recent research integrates physics into DyNVS [26, 31], it largely remains limited to single materials and often single objects within scenes. As a result, the crucial aspect of physics reasoning for interactions between multiple objects in DyNVS—particularly for the estimation of force exchange and deformation during contact—remains largely unexplored. We believe our novel dataset—specifically designed to capture complex multi-body interactions—will serve as a valuable resource for advancing future research in multi-physics modeling and adaptive representations for hybrid scenes.
Integrating material-specific physics solvers
@@
The dominant approach to incorporating physics into DyNVS algorithms currently relies on differentiable particle-based simulators [48, 65, 66, 67, 68, 69], which treat Gaussian primitives as particles. However, different physical phenomena are best captured by different physics solvers—a principle reflected in our dataset construction process described in Section 4. This insight can guide researchers seeking to integrate more appropriate solvers tailored to specific material behaviors, such as fluids, cloth, and smoke. For instance, the FLIP [70] solver excels in simulating incompressible fluids due to its hybrid particle-grid representation, offering improved stability and realism compared to purely particle-based SPH-based solvers. Similarly, thermodynamic phenomena, such as smoke dynamics, are conventionally handled by voxel-based simulators. Nevertheless, the seamless integration of these volumetric solvers into particle-based representations like Gaussian Splatting presents a significant and largely unexplored research challenge.
4 Dataset Construction
Our dataset consists of 17 scenes, categorized into four material types: liquids, gases, viscoelastic materials, and textiles. Each category contains 4 to 5 scenes. This section details the solver properties for each material and provides key simulation details regarding multi-body interactions. Please refer to our supplemental document for other simulation configurations.
4.1 Common Setup
Simulator selection
Rendering
All frames are rendered at a resolution of $640\times 720$. We use path tracing with 256 samples per pixel and apply NVIDIA OptiX for denoising. Scenes are illuminated with 1–3 point lights (intensity: 600–4000); shadows are disabled in textile-focused scenes like tube-flag to emphasize geometry. All cameras follow spiral trajectories to capture diverse views over time.
4.2 Liquid
For liquid scenes, we adopt the Fluid-Implicit Particle (FLIP) solver [70], a hybrid particle-grid method. FLIP maintains particle velocities throughout the simulation, using the grid solely to compute and apply forces such as pressure and viscosity. This approach preserves fine-scale, high-frequency particle velocities, which is crucial for modeling realistic and rapid liquid behavior while adhering to the Navier–Stokes equations. While the Material Point Method (MPM) [48] can also model fluids, its direct velocity aggregation onto the grid limits its ability to capture highly dynamic fluid phenomena like splashing. Particle-only solvers [71] are another option, but FLIP is generally better suited for incompressible fluids due thanks to its hybrid grid-based representation.
Dynamic interactions
4.3 Gas
Dynamic interactions
4.4 Viscoelastic Substances
We use MPM [48] for viscoelastic substances like snow and jelly. As an extension of FLIP [73], MPM is ideal for simulating chunk-based materials. It aggregates particle information onto a grid, performs computations, and then reprojects results back to particles, effectively capturing deformation and internal force propagation.
Dynamic interactions
4.5 Textile
Dynamic interactions
To simulate interactions between objects and textiles–where both move and exchange forces, as seen particularly in the basin scene–we employ the shape match constraint. This constraint helps maintain the overall shape of objects by driving points toward their rest configuration, allowing the material to preserve its structural integrity while still interacting dynamically with textiles and other objects. For the lucy scene, we increased the simulation sub-step count fivefold over the default to robustly handle collisions with the statue’s complex geometry. To simulate wind-induced fluttering in the flags, single-flag, and tube scenes, we applied external forces using a POP Wind node that blows parallel to the ground plane.
5 Analysis
5.1 Evaluation of Existing Algorithms on PhysGaia
Experimental setup
We evaluate recent DyNVS methods on our PhysGaia dataset: D-3DGS [36], 4DGS [37], STG [38], and SOM [15]. Except for SOM [15], which is specialized for monocular setups, all models were tested under both monocular and multiview setups. For point cloud initialization, we first run COLMAP for dense matching and fusion, then uniformly downsample its output to approximately 40,000 points. All experiments are conducted on a NVIDIA A6000 GPU. For evaluation, we adopt standard image quality metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [77], and Learned Perceptual Image Patch Similarity (LPIPS) [78].
Quantitative results
@@ 5.1 Evaluation of Existing Algorithms on PhysGaia
Table 3: Average quantitative results for each material category across all baseline algorithms. Performance tends to be higher for textile materials, but it notably worsens for viscoelastic materials, which typically exhibit highly complex motions involving numerous independent objects.
Capture Type Method Liquid Gas
Monocular D-3DGS [36] 22.7 0.87 0.22 21.9 0.89 0.16
4DGS [37] 24.2 0.87 0.23 21.7 0.88 0.17
STG [38] 19.2 0.72 0.39 21.9 0.85 0.24
SOM [15] 19.6 0.80 0.32 20.0 0.84 0.27
Multiview D-3DGS [36] 22.2 0.87 0.24 23.7 0.91 0.13
4DGS [37] 25.1 0.88 0.22 24.2 0.89 0.17
STG [38] 20.8 0.75 0.40 25.0 0.91 0.19
Capture Type Method Viscoelastic materials Textile
Monocular D-3DGS [36] 20.1 0.84 0.15 22.1 0.83 0.18
4DGS [37] 19.5 0.82 0.18 24.9 0.84 0.18
STG [38] 13.6 0.63 0.40 21.9 0.84 0.21
SOM [15] 16.7 0.75 0.23 20.7 0.79 0.22
Multiview D-3DGS [36] 22.2 0.89 0.10 27.7 0.90 0.12
4DGS [37] 21.0 0.85 0.15 26.6 0.87 0.15
STG [38] 17.2 0.70 0.36 21.1 0.81 0.25
Tables 3 and 4 report average performance by material type and across all 17 scenes. As expected, multiview setups outperform monocular ones, yet even they often yield PSNR below 30. This underscores the difficulty of capturing complex multi-body dynamics in our dataset, suggesting physics-aware priors are essential beyond RGB-based fitting alone. Among materials, textiles are generally reconstructed more accurately, while viscoelastic materials pose greater difficulty due to their complex, object-rich motion. Interestingly, SOM [15], despite strong results on other datasets, performs inconsistently on ours. We presume its motion factorization struggles with highly dynamic phenomena such as splashing and multiple interacting objects. Additionally, its heavy reliance on external modules like point trackers leads to performance drops when these modules fail in challenging scenes, particularly those lacking distinct visual features like smoke or gas.
Qualitative results
@@
(e) SOM [15]
Figure 5: Qualitative results of recent DyNVS methods on the jelly party, lucy, and hanok scenes with monocular setup. All methods struggle to accurately capture multi-body interactions by frequently exhibiting needle-like artifacts and failing to reconstruct dynamic elements accurately.
Figure 5 presents qualitative results from monocular settings, revealing that all methods struggle to accurately capture multi-body interactions. This leads to common issues like needle-like artifacts and under-reconstruction of dynamic regions. For 4DGS [37], its default grid learning rate often causes NaN loss values; while reducing it stabilizes training, the output resembles static scenes due to poor dynamic capture. More results are available in the supplementary material.
5.2 Limitation in Physical Realism
We analyze a key limitation of recent DyNVS methods regarding physical realism. Since Gaussian Splatting-based representation, used in recent algorithms, is an explicit representation, the motion of gaussian primitives is expected to closely follow the actual physical trajectories of dynamic scenes. However, all tested DyNVS algorithms consistently fail to accurately reconstruct these physically reliable motions, as shown in Figure 6. D-3DGS [36], 4DGS [37], and STG [38] produce only localized fluctuations, while SOM [15] exhibits erratic motion primarily due to inaccuracies from external point trackers. This discrepancy is acutely apparent in fluid or gas scenes. Here, photorealism is easily achieved via color similarity alone, allowing primitives to simply fluctuate locally near the surface rather than genuinely modeling true dynamics. This analysis highlights a critical limitation—relying solely on RGB reconstruction loss ensures photorealism but does not guarantee physical realism. We believe that our PhysGaia dataset, which provides ground-truth trajectories, will be instrumental in advancing future research toward physics-aware dynamic scene reconstruction.
Table 4: Average quantitative results across all 17 scenes for monocular and multiview settings. Multiview setups generally offer better reconstruction performance than monocular ones, yet PSNR scores remain below 30 even with multiple views. This highlights the inherent difficulty of reconstructing complex multi-body interactions within our dataset.
@@ 5.2 Limitation in Physical Realism
(e) SOM [15]
Figure 6: Comparison of reconstructed trajectories and its ground truth on the Box-smoke and Pisa scenes. The reconstructed trajectories do not accurately follow the true physical trajectories.

### v2

@@ 3.1 Benchmark Properties
In contrast to real-world video datasets, where the ease of capture is offset by inaccessible underlying physics, our simulated dataset offers complete access to all physical information. This includes 3D particle trajectories and physics parameters such as viscosity, Young’s modulus, Poisson’s ratio, and temperature for gas scenes [^1] Temperature is a crucial component when adopting smoke-related simulators for physical reasoning, particularly for buoyancy calculation.. This comprehensive provision enables precise evaluation of physical reasoning in dynamic scenes, directly facilitating the future research directions outlined in Section 3.2.
@@
Our benchmark uniquely supports both multi-view and monocular dynamic novel view synthesis (DyNVS) tasks, as shown in Table 1. Unlike most existing multi-view datasets [40, 2, 39, 73] with videos captured from fixed camera positions, PhysGaia provides several moving monocular video sequences with independent trajectories. This allows for diverse training configurations; using full sequences corresponds to a multiview DyNVS task, while using a single sequence corresponds to a monocular DyNVS task. For the evaluation, we use two static cameras with a large baseline.
@@
Figure 3: Examples of diverse modalities that users can generate from the provided simulation node graphs. This can facilitate adaptation to specific downstream tasks.
We provide complete source files encompassing physics solvers, source geometries, materials, and texture controls. Such an open and comprehensive framework enables researchers to generate customized datasets with higher resolution imagery and diverse modalities including depth maps, surface normals, and relighted scenes, facilitating adaptation to specific downstream tasks, as shown in Figure 3.
@@
Our PhysGaia benchmark is designed for research-friendly and easy access, providing integration pipelines that enable the use of state-of-the-art 4D Gaussian Splatting models [70, 79, 74, 42, 38] with our data. Furthermore, we include COLMAP-reconstructed point clouds for each scene. These provisions aim to facilitate the adoption of our dataset by researchers working with state-of-the-art DyNVS models.
@@ 3.2 Potential Research
This subsection highlights the potential impact of our PhysGaia dataset by outlining several promising future research directions it uniquely enables.
Physical reasoning of dynamic scenes
@@
Since our PhysGaia dataset provides ground-truth physics information, it facilitates precise evaluation of physical reasoning in dynamic scenes. For instance, ground-truth physics parameters like viscosity can be used to evaluate inverse physics estimation methods, where differentiable simulators are employed to optimize these parameters. Furthermore, unlike existing 4DGS research that primarily focuses on photorealism and relies on ground-truth RGB images, our dataset offers ground-truth 3D trajectories, enabling evaluation of the actual motion of individual Gaussian primitives in 4DGS. We believe this unique feature establishes our dataset as a valuable benchmark for developing and evaluating physics-aware DyNVS models.
Multi-body interaction
@@
While recent research integrates physics into DyNVS [29, 85, 6, 10], it largely remains limited to single materials and often single objects within scenes. As a result, the crucial aspect of physics reasoning for interactions between multiple objects in DyNVS – particularly the estimation of force exchange and deformation during contact – remains largely unexplored. We believe our novel dataset, specifically designed with complex multi-object interactions, will be instrumental in enabling significant future research in multi-physics modeling and adaptive representations for handling hybrid scenes.
Integration of material-specific physics solver
@@
The dominant approach for integrating physics into DyNVS algorithms currently involves adopting differentiable simulators [66, 26, 51, 1, 24, 25], treating Gaussian primitives as particles in simulators. However, we emphasize that different physical phenomena are best captured by different physics solvers, a principle reflected in our dataset construction process detailed in our supplementary documents. This can guide researchers seeking to integrate more appropriate solvers tailored to specific material behaviors, such as fluids, cloth, or smoke. For example, the FLIP [5] solver excels at simulating incompressible fluids due to its hybrid particle-grid representation, offering greater stability and realism compared to purely particle-based SPH-based solver. Similarly, thermodynamic effects like temperature and buoyancy are crucial for smoke simulation, typically represented using voxel-based grids. Integrating such volumetric solvers into particle-based frameworks like Gaussian splatting, however, remains largely unexplored.
4 Metrics for Physical Realism
When reconstructing 4D scenes, faithfully capturing actual 3D particle flows can be viewed as a proxy for physically realistic reconstruction. To this end, we introduce metrics that quantify how well the motions of Gaussian primitives in 4DGS align with ground-truth physical trajectories.
Trajectory Distance (TD)
Trajectory distance measures the spatial deviation along entire trajectories. Let $M$ denote the number of reconstructed primitives, $T$ the total number of frames, $X_{i}^{t,\mathrm{recon}}$ the 3D position of reconstructed primitive $i$ at time $t$, and $X_{j}^{t,\mathrm{gt}}$ the 3D position of ground-truth trajectory $j$. TD is defined as the average Euclidean distance between matched trajectories as follows:
$\mathrm{TD}=\frac{1}{MT}\sum_{i=0}^{M-1}\sum_{t=0}^{T-1}\|X_{i}^{t,\mathrm{recon}}-X_{j(i)}^{t,\mathrm{gt}}\|_{2},$ (1)
where $j(i)$ denotes the closest ground-truth trajectory matched to primitive $i$ in the initial frame. Lower TD values indicate closer alignment with ground-truth motion, reflecting higher physical realism.
Area Under the Outlier Percentages (AUOP)
To complement TD, we introduce a percentage-based metric that captures outlier behavior holistically across the entire sequence. For each primitive $i$ at timestep $t$, an outlier indicator $O_{i}^{t}$ is defined as follows:
$O_{i}^{t}=\begin{cases}1&\text{if }O_{i}^{t-1}=1\text{ or }\|X_{i}^{t,\mathrm{recon}}-X_{j(i)}^{t,\mathrm{gt}}\|_{2}>\delta,\\ 0&\text{otherwise},\end{cases}$ (2)
where $1$ indicates the primitive is marked as an outlier, and $\delta$ is a deviation threshold. Once a trajectory exceeds this threshold, it remains marked as an outlier for all subsequent frames. The percentage of outlier trajectories is computed at each timestep, yielding a curve over time as shown in Figure 8. To summarize this temporal behavior, we propose AUOP, which quantifies the area under this curve and provides a single comprehensive value for the entire sequence. Lower AUOP values indicate better compliance with ground-truth physics, reflecting fewer and shorter deviations throughout the sequence.
Discussion
TD and AUOP provide complementary perspectives: TD measures absolute trajectory accuracy, while AUOP captures the persistence of large deviations, together enabling a more robust evaluation of physical realism in reconstructed flows.
5 Empirical Analysis with PhysGaia
In this section, we demonstrate the characteristics of our benchmark and highlight critical yet largely overlooked limitations of existing DyNVS approaches.
5.1 Comparison with Physics-based Benchmarks
Table 3: Comparison with existing physics-based benchmarks on motion complexity and photorealism. PhysGaia achieves the highest dynamic score [44] and lowest FID [23] and KID [3], demonstrating richer dynamics and greater visual realism than existing DyNVS datasets. Note that PAC-NeRF sequences are too short to compute a dynamic score.
Benchmark Motion complexity Photorealism
Multi-body Dynamic score $\uparrow$ FID $\downarrow$ KID $\downarrow$
ScalarFlow [14] No 0.391 293.5 0.255
PAC-NeRF [41] No N/A 242.6 0.164
Spring-Gaus [85] No 0.372 261.8 0.171
PhysGaia Yes 0.444 207.8 0.118
(a) ScalarFlow [14]
(b) PAC-NeRF [41]
(c) Spring-Gaus [85]
Figure 4: Limitations of existing datasets. While ScalarFlow, PAC-NeRF, and Spring-Gaus address physical phenomena, they are limited in narrow coverage of physical materials, overly simplified dynamics, and an absence of rich multi-object interactions.
Existing physics-based benchmarks [14, 41, 85] [^2] These datasets provide reliable ground-truth physics, including trajectories and physical parameters, as detailed in Table 2. often exhibit limited motion diversity and simplistic dynamics, as visualized in Figure 4. Table 3 demonstrates that PhysGaia significantly bridges the gap in both motion complexity and realism. To evaluate motion complexity, we employ the dynamic score [44], which quantifies motion through feature-level distances and inter-frame variance. PhysGaia achieves the highest score, reflecting its capacity to capture intricate multi-body interactions. Regarding photorealism, we compute FID [23] and KID [3] by flattening video sequences into image sets to measure distributional alignment with TinyImageNet. Our dataset yields the lowest scores across both metrics, confirming superior visual fidelity. These results validate that PhysGaia effectively bridges the gap in existing benchmarks by integrating physical accuracy, high photorealism, and complex motion dynamics.
5.2 Analysis on Dynamic Novel View Synthesis
We analyze recent DyNVS approaches (D-3DGS [79], 4DGS [74], STG [42], MoSca [38], and SoM [70]) on PhysGaia in terms of both photorealism and physical realism.
5.2.1 Evaluation of Photorealism
@@
Table 4: Average quantitative results for each material category across all algorithms. While performance is generally high for textile, it deteriorates significantly for rheological substances, which typically exhibit complex dynamics involving multiple interacting components.
View Method Liquid Gas Rheological materials Textile
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Monocular D-3DGS [79] 22.7 0.87 0.22 21.9 0.89 0.16 20.1 0.84 0.15 22.1 0.83 0.18
4DGS [74] 24.2 0.87 0.23 21.7 0.88 0.17 19.5 0.82 0.18 24.9 0.84 0.18
STG [42] 19.2 0.72 0.39 21.9 0.85 0.24 13.6 0.63 0.40 21.9 0.84 0.21
MoSca [38] 20.5 0.78 0.39 21.2 0.79 0.35 17.8 0.74 0.23 18.6 0.74 0.36
SoM [70] 19.6 0.80 0.32 20.0 0.84 0.27 16.7 0.75 0.23 20.7 0.79 0.22
Multiview D-3DGS [79] 22.2 0.87 0.24 23.7 0.91 0.13 22.2 0.89 0.10 27.7 0.90 0.12
4DGS [74] 25.1 0.88 0.22 24.2 0.89 0.17 21.0 0.85 0.15 26.6 0.87 0.15
STG [42] 20.8 0.75 0.40 25.0 0.91 0.19 17.2 0.70 0.36 21.1 0.81 0.25
@@
(e) MoSca [38]
(f) SoM [70]
Figure 5: Qualitative results of recent DyNVS methods on the jelly party scene with monocular setup. All methods struggle to accurately capture multi-body interactions by frequently exhibiting needle-like artifacts and failing to reconstruct dynamic elements accurately. Please refer to our supplementary documents for more qualitative results.
Figure 6: Comparison of motion diversity across DyNVS datasets. Dynamic Score (DS) [44] measures feature-level distances across frames, serving as a proxy for motion complexity. PhysGaia achieves the highest score, reflecting the rich multi-object interactions present in our benchmark.
We adopt standard image quality metrics for photorealism evaluation: PSNR, SSIM [72], and LPIPS [82]. Tables 4 and E show that existing methods struggle on PhysGaia, unlike their performance on existing DyNVS benchmarks [80, 18]. We attribute this to the motion complexity from multi-object interactions, evidenced by the significantly higher motion entropy of PhysGaia (Figure 6). Each method faces distinct challenges: MLP- and grid-based representations (D-NeRF [57], 4DGS [74]) fail on fine-grained dynamics such as fluid splashing, polynomial motion models (STG [42]) are too restrictive for complex trajectories, and motion factorization or ARAP-based constraints (MoSca [38], SoM [70]) break down under rich multi-object interactions. Figure E illustrates these failures, where all methods produce needle-like artifacts in the jelly party scene. Additional qualitative results are provided in the supplementary material.
Table 5: Average quantitative results for both monocular and multiview settings, averaged across all 17 scenes. While multiview setups generally offer better reconstruction performance than monocular ones, even multiview results achieve PSNR scores below 30. This highlights the substantial difficulty in reconstructing the complex multi-body interactions in our benchmark.
@@ 5.2.2 Evaluation of Physical Realism
(e) MoSca [38]
(f) SoM [70]
Figure 7: Comparison of reconstructed flows and ground truth on the Box-smoke and Pisa scenes. Reconstructed flows deviate significantly from ground truth, where photorealism is achieved through local surface fluctuations rather than following actual physical motion.
Figure 8: Outlier percentage over time on the box-smoke scene with AUOP values indicated. The oracle (GT-GT) achieves near-zero outliers, while existing methods exhibit rapidly increasing outlier percentages, resulting in significantly higher AUOP values. This highlights their inability to track physically plausible flows.
Table 6: Performance for physics realism. The box-smoke scene shows higher TD and AUOP; existing methods are hard to capture physics with simplified visuals.
Method Cow Box-smoke
TD $\downarrow$ AUOP $\downarrow$ TD $\downarrow$ AUOP $\downarrow$
D-3DGS [79] 0.07   0.0 4.01 37.9
4DGS [74] 0.90   0.0 6.00 68.8
STG [42] 0.30   0.7 3.62 30.2
MoSca [38] 0.33   0.7 3.67 28.3
SoM [70] 0.61 17.7 4.45 25.4
Oracle 0.03   0.0 0.84   0.3
To evaluate whether Gaussian primitives follow ground-truth physical flows, we report TD and AUOP for two scenes in Table 6333Oracle denotes results from an additional set of ground-truth trajectories (GT-GT), serving as an upper bound on performance.: the less dynamic cow scene and the highly dynamic box-smoke scene. While all methods perform reasonably on the cow scene, they exhibit substantially higher TD and AUOP on the box-smoke scene—optimizing solely for RGB reconstruction achieves photorealism but fails to ensure physical plausibility. Figures 7 and 8 further confirm this: reconstructed trajectories deviate significantly from ground truth, with primitives merely fluctuating near the surface rather than following actual physical motion, and outlier percentages increasing rapidly over time. These results underscore the necessity of physics-aware evaluation and position PhysGaia as a critical benchmark for driving progress in this direction.
5.3 Analysis on Inverse-physics Problems
Table 7: Comparison of estimated physics parameters (Young’s modulus $E$ and Poisson’s ratio $\nu$) against ground truth. Both PAC-NeRF [41] and GIC [6] methods substantially underestimate material stiffness in multi-object scenarios.
   Method    $E$    $\nu$
   Ground Truth    $100.0\times 10^{3}$    0.50
   PAC-NeRF [41]    $1.7\times 10^{3}$    0.42
   GIC [6]    $23.4\times 10^{3}$    0.37
Beyond DyNVS tasks, PhysGaia also supports physics parameter estimation [6, 85, 10] since it is faithfully generated using physics simulators with known ground-truth parameters. We evaluate both PAC-NeRF [41] and GIC [6] on the bouncing balls scene. As shown in Table 7, both methods exhibit similar trends and significantly underestimate Young’s modulus $E$. This behavior arises because both approaches were designed for single-object scenarios: when applied to multi-object scenes with complex motions and interactions, the algorithms favor lower $E$ values, allowing greater deformation to fit the observations. These results demonstrate that our benchmark exposes underexplored challenges in multi-object physics parameter estimation, providing a valuable testbed for advancing this research direction.